# Knowledge Augmentation in Federation: Rethinking What Collaborative Learning Can Bring Back to Decentralized Data

### v2

@@ III-C Outline of Approaches
Through in-depth investigation, we figure out that there has been early effort in related areas especially in the conjunction of FL and data engineering. Even though the motivation behind these studies do not explicitly align with the goal of KAF, we found their methodologies somewhat suitable for serving as the building blocks for KAF. By categorizing these studies, we present a suggested roadmap for KAF comprised of three approaches: federated knowledge expanding, federated knowledge filtering, and federated label and feature space correction (Fig. 3.
Figure 3: The structure of this article.
IV Methods for Knowledge Expansion
As discussed previously, a federated system entails a set of knowledge domains that collectively comprise the global domain. Intuitively, it is easy to understand that greater overlaps between the local domains can facilitate faster and better learning. For example, it is commonly agreed that the global model’s convergence can be improved by reducing the “non-IIDness” of local data. In fact, this can be achieved by outspreading the knowledge contained in each local domain.
In this section, we investigate popular approaches for collaborative knowledge expansion under the FL setting, where the implications of expansion are two-fold. First, this represents a roadmap for extending the boundaries of local knowledge domains through introducing exogenous information from other clients. Second, the expansion of knowledge also refers to the augmentation of data with the aims of re-balancing local distribution and generating diversified instances. Hence, the key criteria for this research category are:
  • Seeking to extend local domains of knowledge through collaboration;
  • Improving the diversity of data samples within each local domain based on information exchange.
@@ IV Methods for Knowledge Expansion
We summarize that there are two major lines of research in this direction. In the following content, we first introduce existing studies for extrapolating the knowledge in local data by information sharing and reconstruction in a federated system. Then we revisit the research on enhancing local data in terms of size, quality and distribution via data augmentation.
IV-A Federated Knowledge Extrapolation
@@
We conceptually describe knowledge extrapolation as the methodology of filling the missing part of knowledge that is not explicitly expressed by the raw data. This process usually involves predicting or approximating implicit samples locally by utilizing some global knowledge contributed and shared by other participants in the system (Fig. 4). In the context of federated learning, knowledge extrapolation can be very useful when all or most of the local datasets are sparse in terms of information density. For example, Sanyal et al. [42] considered a healthcare scenario where numerous Internet of Medical Things (IoMT) devices are simultaneously working towards sensing a particular phenomenon. In this case each IoMT device acts as an independent data collector that simultaneously generate time series data. The challenge here is that these readings need to be collectively utilized for analysis, but it is not encouraged to continuously push them to a central server due to privacy concerns and network constraints. To this end, the authors proposed a Federated Filtering Framework (FFF) [42], where the key component on each local device is a predictor for approximating the true sensing data. The IoMT devices communicate with a fog server that aggregates and utilizes the predictors in order to reconstruct the full data matrix. For communication efficiency, the timing of local model update and aggregation is determined by dynamic thresholds.
@@
Figure 4: A schematic showing the general framework of knowledge extrapolation. The goal is to obtain the missing part of knowledge in the local domain.
@@
Sparsity is one of the major obstacles that can largely undermine the efficacy of machine learning. The problem has been extensively explored in the context of recommender systems and the design of recommendation algorithms. More specifically, data are represented as a user-item interaction matrix that is huge in dimensionality but low in information density, which is caused by the natural pattern of user behavior and the scale of e-commerce platforms. Although there are many solutions to it in the literature, most of the existing studies make the assumption that the whole data matrix is available for analysis. However, a centralized collection of user behavior data may not be permitted in many situations such as cross-border e-commerce.
@@
As one of the most popular approaches to recommender systems, Collaborative Filtering (CF) an be viewed as a data model that is able to extrapolate incomplete knowledge (i.e., unrecorded interactions) using the dense representation of factor matrices. Approved by the Netflix Prize, Matrix Factorization (MF) is the classic implementation of CF, where the raw user-item interaction matrix $\boldsymbol{R}\in\mathbb{R}^{m\times n}$ can be approximated as
@@
By adapting MF to FL settings, researchers have found new paths to personalized recommendation without a centralized storage of user vectors (which are likely to contain user-private information). Ammad-ud-din et al. [43] first described this new approach as Federated Collaborative Filtering (FCF), where all the clients keep their interaction data private on their own devices. Based on a reconstruction loss and stochastic gradient descent (SGD), clients in FCF update their user-specific (user-factor) matrices locally given the latest item-factor matrix from the server. In the meantime, each client also calculates its corresponding part of the item-specific gradients and feeds back to the server for item-factor matrix update. Following a similar framework, Yang et al. [44] further investigated a multi-behavior scenario, where both item viewing and purchase are recorded but the former is kept private by each user. The authors proposed to binarize both factor matrices and utilized a caching mechanism to reduce communication payload for collaborative parameter update. Considering that a single form of information may not be sufficient for accurate recommendation, the knowledge provided by implicit and explicit feedback can be integrated using a probabilistic model. In this case, a specific decomposing strategy is needed for model updating without exposing users’ private information [45].
Neural Collaborative Filtering (NCF) [46] replaces the matrix factorization process of CF by using deep neural networks (DNNs) to learn the user embedding vectors $\boldsymbol{X}=\{\boldsymbol{x_{1}},\boldsymbol{x_{2}},\ldots\}$ and item embedding vectors $\boldsymbol{Y}=\{\boldsymbol{y_{1}},\boldsymbol{y_{2}},\ldots\}$. With an embedding dimensionality of $k$, these learned latent vectors are then concatenated and fed into a Multi-Layer Perceptron (MLP) that outputs the predicted score of preference.
@@
NCF differs from MF-based approaches in the sense that knowledge is encapsulated in deep neural networks (DNNs), rather than factor matrices. In such way each user and item can be encoded independently and fed into a prediction model (e.g., another DNN) for training and inference. Inspired by this property, Perifanis and Efraimidis [47] designed a federated NCF framework that extends that basic NeuMF model where each client holds three sets of parameters including their user profile, item profile, and the DNN’s weights. With user profiles kept private, they proposed a secure aggregation algorithm, MF-SecAgg, to perform server-side parameter update, with separate steps for item profile update and DNN update respectively. With a similar motivation, Jiang et al. [48] introduced an adaptive Differential Privacy (DP) mechanism with decaying noise into the communication process of Federated NCF. This protects the exchange of model gradients against inference attacks.
@@
For effective knowledge extrapolation, it is intricate to strike a balance between preserving learned representations locally and leveraging the common knowledge shared globally. To address this, Li et al. [49] designed a gating dual-encoder structure based on Variational Auto-Encoder (VAE) for personalized CF in a federated setting. They introduced a gating network followed by two encoders, one for pertaining local knowledge and the other for learning user representations in the global latent space. By mapping the user-item interaction vectors to separate latent subspaces, the loss of personalized knowledge (due to federated aggregation) is effectively mitigated.
@@
Generally, user-private knowledge needs to be stored locally, no matter as a factor matrix (for MF-based methods) or a neural network’s parameters (for NCF). This facilitates global knowledge fusion but also increases the burden on end devices. One may ask that is it possible to make local knowledge “invisible”? Singhal et al. [50] offered their answer by presenting a novel reconstruction-based framework FedRecon. Inspired by Model-Agnostic Meta Learning (MAML), FedRecon allows stateless participants to rebuild their local model every time they join the federation with the knowledge provided by the global model. This facilitates more flexible collaboration at scale and, more importantly, indicates that personalized knowledge can be extrapolated by leveraging globally shared knowledge.
@@ IV-B Federated Data Augmentation
Data augmentation, as a powerful technique for diversifying the data, can also enable effective expansion of knowledge when deployed in a federated system. The value of federated data augmentation is two-fold. On the one hand, it directly introduces new knowledge—typically represented by synthetic samples—to the set of local data. On the other hand, multiple data owners can collaboratively perform data augmentation under the guidance of a global coordinator, which brings higher consistency of data distribution across the parties in the federation.
@@
The implementations of augmentation generally involve applying transforms on existing samples or synthesizing new samples based on a priori. Note that the synthesis of data and the training of downstream models can be decoupled, hence a simple implementation is to perform data augmentation in the local pre-processing stage without interleaving with the following FL process. Kundu et al. [51] followed this approach and developed a federated monkeypox disease detection model on the basis of offline image augmentation. Specifically, a Cycle GAN is pre-trained and deployed locally to generate additional samples to enhance local training. Similarly, Ullah et al. [52] developed an FL-based smart healthcare system, where they employed oversampling in order to address the class imbalance of chest X-ray imaging data. More specifically, SMOTE is adopted to mitigate the scarcity of minority-class samples such as those with “COVID-19” and “Viral” labels [53]. One potential limit is that unreasonable samples could be generated without using domain-specific knowledge as constraints. Despite being easy-to-implement, offline data augmentation is agnostic of any global knowledge and thus may not serve well for the global model’s convergence.
The potential limit of offline data augmentation can be explained by considering the boundary of knowledge for each domain, wherein offline data augmentation is essentially analogous to linear and non-linear interpolation within the domain. By contrast, through introducing exogenous information for local data augmentation (i.e., collaborative data augmentation), more generalizable knowledge can be obtained to expand the domain. This could lead to higher gains for the downstream tasks.
@@
TABLE II: Summarizing the studies related to federated knowledge expansion in terms of methodological highlights, drawbacks, datasets for experiments, and the form of knowledge for exchange.
@@
[42] Fed. knowledge extrapolation predictors & filters $\bullet$ remote data reconstruction via predictors
@@
[43, 45] Fed. knowledge extrapolation factor matrices $\bullet$ decoupling private and shareable knowledge via matrix factorization $\bullet$ linear growth of communication cost with the raw matrix’s dimensions MovieLens, FilmTrust, Epinions, and in-house data
[44] Fed. knowledge extrapolation binarized matrices $\bullet$ finer categorization of private knowledge
@@
[47, 48, 49] Fed. knowledge extrapolation latent vectors and prediction models $\bullet$ better protection of private knowledge
@@
[50] Fed. knowledge extrapolation the global model $\bullet$ model-agnostic and stateless collaboration
@@
[51, 52] Federated data augmentation task-specific models $\bullet$ easy to implement
@@
[54] Federated data augmentation attention scores from generative models $\bullet$ CVAE for local augmentation
@@
[55, 56] Federated data augmentation the average of local samples $\bullet$ controllable anonymity
@@
[57, 58] Federated data augmentation prediction model $\bullet$ Augmented images with cross-institute knowledge $\bullet$ tightly-coupled with the task fastMRI and IXI
[59] Federated data augmentation shared generator $\bullet$ decoupled knowledge with decoupled models $\bullet$ imbalanced amount of knowledge for the generator and discriminator fastMRI, IXI, BRATS, and in-house data
Although the collaboration of different clients naturally provides a rich set of knowledge sources for data augmentation, how to share the knowledge without compromising user privacy remains the key challenge. Essentially, a client needs to perceive the distribution of data outside its domain without raw data exchange. On this point, sharing a generator (i.e., a generative model that produces new samples) has proved to be an useful mechanism. Wen et al. [54] proposed a generative models-based federated data augmentation strategy (FedDA), where a Conditional AutoEncoder (CVAE) is deployed and collaboratively trained on each client to tackle non-IID data distribution. Instead of sharing the whole generator, they extracted the decoder’s hidden-layer features as the shared knowledge and introduced a knowledge distillation loss to guide the training. This effectively reduces communication overheads. For privacy preservation, noise samples are mixed into each local dataset. Intuitively, one can adapt traditional data augmentation, such as the popular MixUp [38] strategy, to the federated setting. However, the difficulty here is that raw samples need to be shared if we want to mix up local and global knowledge. Over the generated samples $(\boldsymbol{x}^{\prime},y^{\prime})$ this “global MixUp” strategy theoretically corresponds to minimizing the following loss:
$\ell_{rawmix}(h_{\theta}(\boldsymbol{x}^{\prime}),y^{\prime})=\ell\Big{(}h_{\theta}\big{(}(1-\lambda)\boldsymbol{x}_{i}+\lambda\boldsymbol{x}_{j}\big{)},(1-\lambda)y_{i}+\lambda y_{j}\Big{)}$ (4)
where $\ell$ and $h_{\theta}$ is the task-specific loss criterion and the model to be trained.
To achieve privacy preservation, Yoon et al. [55] first presented Mean Augmented Federated Learning (MAFL), a simple framework that enables clients to exchange the average of local samples (called “mashed” data) along with model updates. Further, through approximating the global MixUp loss (Eq. 4) with Taylor expansion, they derived that the approximation can be done using local samples and the averaged samples from other clients (Eq. (5)):
This motivated them to devise the FedMix [55] algorithm that performs collaborative MixUp augmentation and training based on a decomposable loss function. Zhang et al. [56] further extended this framework for regression tasks using bilateral neighborhood expansion and evaluated it on various public datasets. Empirical results show that their algorithms yielded stronger performance than FedMix [55].
A special case of federated data augmentation is the reconstruction of samples using knowledge from multiple domains, which has proved very useful for accelerated medical imaging such as Magnetic Resonance Imaging (MRI) [57, 58]. To adapt to the disparity of imaging operators between different sites, Elmas et al. [59] deployed an unconditional GAN at each site to generate synthetic high-quality samples, with a private discriminator to fit site-specific knowledge and a globally shared generator for prior knowledge embedding. A tailored reconstruction loss was used to guide the training of local generators.

### v3

@@ III-C Outline of Approaches
Through an in-depth investigation, we collect a rich set of studies that align or partially align with the characterized scenario of TAF. Some of these papers rest in the scope of federated learning but employ data-centric methods suitable for task-agnostic scenarios. We adopt a two-layer taxonomy: in the top layer we define three types of data-centric collaboration: collaborative data expansion, collaborative data refinement, and collective harmonization of data, and in the bottom layer we group the papers into multiple branches based on their technical design. Fig. 3 visualizes the taxonomy of studies covered by this survey.
IV Collaborative Data Expansion
In this section, we investigate popular approaches for collaborative data expansion under the federated setting. By expansion we refer to methods that outspread the knowledge contained in each local domain by diversifying data samples.
@@ IV Collaborative Data Expansion
We summarize that there are two major lines of research in this direction. In the following content, we first introduce existing studies for extrapolating the samples in local data by information sharing and reconstruction. Then we revisit the research on enhancing local data in terms of size, quality and distribution via data augmentation.
IV-A Federated Sample Extrapolation
@@
We conceptually describe sample extrapolation as the methodology of filling the missing part of information that is not explicitly expressed by the raw data. This process usually involves predicting or approximating implicit samples locally by utilizing some global knowledge contributed and shared by other participants in the system (Fig. 4). In a federation, sample extrapolation can be very useful when all or most of the local datasets are sparse in terms of information density. For example, Sanyal et al. [Sanyal2019fff] considered a healthcare scenario where numerous Internet of Medical Things (IoMT) devices act simultaneously as independent data collectors. The challenge here is that their readings need to be collectively utilized for analysis, but it is not encouraged to continuously push them to a central server. To this end, the authors proposed a Federated Filtering Framework (FFF) [Sanyal2019fff], where the key component on each local device is a predictor for approximating the true sensing data. The IoMT devices communicate with a fog server that aggregates and utilizes the predictors in order to reconstruct the full data matrix.
@@
Figure 4: A schematic showing the general framework of federated sample extrapolation. The goal is to obtain the missing part of knowledge in the local domain.
@@
Sparsity is one of the major obstacles that can largely undermine the efficacy of machine learning and data mining. The problem has been extensively explored in the context of recommender systems where the data are represented as a user-item interaction matrix that is huge in dimensionality but low in information density, which is caused by the natural pattern of user behavior and the scale of e-commerce platforms. Although there are many solutions to it in the literature, most of the existing studies make the assumption that the whole data matrix is available for analysis. However, a centralized collection of user behavior data may not be permitted in many situations such as cross-border e-commerce.
@@
As one of the most popular approaches to recommender systems, Collaborative Filtering (CF) can be viewed as a data model that is able to extrapolate incomplete information (i.e., unrecorded interactions) using the dense representation of factor matrices. Approved by the Netflix Prize, Matrix Factorization (MF) is the classic implementation of CF, where the raw user-item interaction matrix $\boldsymbol{R}\in\mathbb{R}^{m\times n}$ can be approximated as
@@
It is hard to predict unknown interactions (i.e., new knowledge) without a centralized storage of user vectors (which are likely to contain user-private information). Ammad-ud-din et al. [FCF2019] first described this new model as Federated Collaborative Filtering (FCF), where all the participants keep their interaction data private. Based on a reconstruction loss and stochastic gradient descent (SGD), participants in FCF update their user-specific (user-factor) matrices locally given the latest item-factor matrix from the server. In the meantime, each participant also calculates its corresponding part of the item-specific gradients and feeds back to the server for item-factor matrix update. Following a similar framework, Yang et al. [yang2024discrete] further investigated a multi-behavior scenario, where both item viewing and purchase are recorded but the former is kept private by each user. The authors proposed to binarize both factor matrices and utilized a caching mechanism to reduce communication payload for collaborative parameter update. Considering that a single form of information may not be sufficient for accurate recommendation, the information provided by implicit and explicit feedback can be integrated using a probabilistic model [yang2022dpmf].
@@
Neural Collaborative Filtering (NCF) [NCF2017] replaces the matrix factorization process of CF by using deep neural networks (DNNs) to learn the user embedding vectors $\boldsymbol{X}=\{\boldsymbol{x_{1}},\boldsymbol{x_{2}},\ldots\}$ and item embedding vectors $\boldsymbol{Y}=\{\boldsymbol{y_{1}},\boldsymbol{y_{2}},\ldots\}$. NCF differs from MF-based approaches in the sense that knowledge is encapsulated in deep neural networks (DNNs) rather than factor matrices. Inspired by this property, Perifanis and Efraimidis [perifanis2022federated] designed a federated NCF framework that extends the basic NeuMF model, where each participant holds three sets of parameters including their user profile, item profile, and the DNN’s weights. With user profiles kept private, they proposed a secure aggregation algorithm, MF-SecAgg, to perform server-side parameter update, with separate steps for item profile update and DNN update respectively. With a similar motivation, Jiang et al. [JiangFedNCF2022] introduced an adaptive Differential Privacy (DP) mechanism with decaying noise into the communication process of Federated NCF. This protects the exchange of gradients against inference attacks.
@@
For effective sample extrapolation, it is intricate to strike a balance between preserving learned representations locally and leveraging the common knowledge shared globally. To address this, Li et al. [li2024personalized] designed a gating dual-encoder structure based on Variational Auto-Encoder (VAE) for federated CF. They introduced a gating network followed by two encoders, one for pertaining local knowledge and the other for learning user representations in the global latent space. By mapping the user-item interaction vectors to separate latent subspaces, the loss of personalized knowledge (due to federated aggregation) is effectively mitigated.
@@
Generally, user-private knowledge needs to be stored locally, no matter as a factor matrix (for MF-based methods) or a neural network’s parameters (for NCF). This facilitates global knowledge fusion but also increases the burden on end devices. So is it possible to make local knowledge “invisible”? Singhal et al. [singhal2021federated] offered their answer by allowing stateless participants to rebuild their local model every time they join the federation with the knowledge provided by the global model. This enables flexible collaboration at scale and, more importantly, indicates that personalized knowledge can be extrapolated by leveraging globally shared knowledge.
@@ IV-B Federated Data Augmentation
Data augmentation is a powerful technique for diversifying the data when deployed in a federated system. The value of federated data augmentation is two-fold. On the one hand, it directly introduces new knowledge—typically represented by synthetic samples—to the set of local data. On the other hand, multiple participants can collaboratively perform data augmentation, which brings higher consistency of data distribution across the parties in the federation.
@@
Data augmentation generally involves applying transforms on existing samples or synthesizing new samples based on a priori. Note that the synthesis of data and the training of downstream models can be decoupled, hence a simple implementation is to perform data augmentation in the local pre-processing stage without interleaving with the following learning process. Kundu et al. [Kundu2024monkeypox] followed this approach and developed a federated monkeypox disease detection model on the basis of offline image augmentation. Specifically, a Cycle GAN is pre-trained and deployed locally to generate additional samples to enhance local training. Similarly, Ullah et al. [Ullah2024xray] developed an FL-based smart healthcare system, where they employed offline oversampling in order to address the class imbalance of chest X-ray imaging data. More specifically, SMOTE is adopted to mitigate the scarcity of minority-class samples such as those with “COVID-19” and “Viral” labels [Chowdhury2020covid19]. Despite being easy-to-implement, offline data augmentation is agnostic of any global knowledge and thus may not serve well for the downstream tasks. The potential limit of offline data augmentation can be explained by considering the boundary of knowledge for each domain, wherein offline data augmentation is essentially analogous to linear and non-linear interpolation within the domain. By contrast, through introducing exogenous information for local data augmentation (i.e., collaborative data augmentation), more generalizable knowledge can be obtained to expand the domain, leading to higher gains for the downstream tasks.
Although the collaboration of different parties naturally provides a rich set of knowledge sources for data augmentation, how to share the knowledge without compromising user privacy remains the key challenge. Essentially, a participant needs to perceive the distribution of data outside its domain without raw data exchange. On this point, sharing a generator (i.e., a generative model that produces new samples) has proved to be an useful mechanism. Wen et al. [Wen2022CVAE] proposed a generative model-based federated data augmentation strategy termed FedDA. They deployed and collaboratively trained a Conditional AutoEncoder (CVAE) on each participant. Instead of sharing the whole generator, they extracted the decoder’s hidden-layer features as the shared knowledge and introduced a knowledge distillation loss to guide the training. This effectively reduces communication overheads. For privacy preservation, noise samples are mixed into each local dataset.
Intuitively, one can adapt traditional augmentation, such as the popular MixUp [Zhang2018mixup] strategy, to the TAF setting. Given a loss criterion $\ell$ and a model $h_{\theta}$, naïve Mixup globally yields the following loss over any generated samples $(\boldsymbol{x}^{\prime},y^{\prime})$:
$\ell_{rawmix}(h_{\theta}(\boldsymbol{x}^{\prime}),y^{\prime})=\ell\Big(h_{\theta}\big((1-\lambda)\boldsymbol{x}_{i}+\lambda\boldsymbol{x}_{j}\big),(1-\lambda)y_{i}+\lambda y_{j}\Big),$ (2)
where the raw data from two participants $i$ and $j$ are involved, posing privacy issues.
To achieve privacy-preserving augmentation, Yoon et al. [yoon2021fedmix] presented Mean Augmented Federated Learning (MAFL), a simple framework that enables participants to exchange the average of local samples (called “mashed” data) along with model updates. Further, through approximating the global MixUp loss (Eq. 2) with Taylor expansion, they derived that the approximation can be done using local samples and the averaged samples from other participants (Eq. (3)):
This motivated them to devise the FedMix [yoon2021fedmix] algorithm that performs collaborative MixUp augmentation and training based on a decomposable loss function. Zhang et al. [Zhang2023fedm] further extended this framework for regression tasks using bilateral neighborhood expansion and evaluated it on various public datasets. Empirical results show that their algorithms yielded stronger performance than FedMix [yoon2021fedmix].
A special case of federated data augmentation is the reconstruction of samples using knowledge from multiple domains, which has proved very useful for accelerated medical imaging [yan2024cross, ahmed2025fedgraphmri]. To adapt to the disparity of imaging operators between different sites, Elmas et al. [elmas2023mri] deployed an unconditional GAN at each site to generate synthetic high-quality samples, with a private discriminator to fit site-specific knowledge and a globally shared generator for prior knowledge embedding.
@@
TABLE II: Summarizing the studies related to collaborative data expansion in terms of methodological highlights, drawbacks, datasets for experiments, and the form of knowledge for exchange.
@@
[Sanyal2019fff] Fed. sample extrapolation predictors & filters $\bullet$ remote data reconstruction via predictors
@@
[FCF2019, yang2022dpmf] Fed. sample extrapolation factor matrices $\bullet$ decoupling private and shareable knowledge via matrix factorization $\bullet$ linear growth of communication cost with the raw matrix’s dimensions MovieLens, FilmTrust, Epinions, and in-house data
[yang2024discrete] Fed. sample extrapolation binarized matrices $\bullet$ finer categorization of private knowledge
@@
[perifanis2022federated, JiangFedNCF2022, li2024personalized] Fed. sample extrapolation latent vectors and prediction models $\bullet$ better protection of private knowledge
@@
[singhal2021federated] Fed. sample extrapolation the global model $\bullet$ model-agnostic and stateless collaboration
@@
[Kundu2024monkeypox, Ullah2024xray] Federated data augmentation task-specific models $\bullet$ easy to implement
@@
[Wen2022CVAE] Federated data augmentation attention scores from generative models $\bullet$ CVAE for local augmentation
@@
[yoon2021fedmix, Zhang2023fedm] Federated data augmentation the average of local samples $\bullet$ controllable anonymity
@@
[yan2024cross, ahmed2025fedgraphmri] Federated data augmentation prediction model $\bullet$ Augmented images with cross-institute knowledge $\bullet$ tightly-coupled with the task fastMRI and IXI
[elmas2023mri] Federated data augmentation shared generator $\bullet$ decoupled knowledge with decoupled models $\bullet$ imbalanced amount of knowledge for the generator and discriminator fastMRI, IXI, BRATS, and in-house data